OpenStudy (anonymous):

l y+3 l +5 = 2y Help :)

hero (hero):

Subtract 5 from both sides, then take the inverse absolute value of both sides.

OpenStudy (anonymous):

when you subtract 5 from both sides you will have |y+3|=2y-5 now the next step is to write the positive and negative y+3= +(2y-5) and y+3= -(2y-5) now solve each for y

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

so you have it now?

OpenStudy (anonymous):

yes I understand it I was confused about subtracting the 5 and negating the other side. but thanks to you..

OpenStudy (anonymous):

Great! just remember that |x| =y you will get that x =y and x=-y

OpenStudy (anonymous):

could you help with one more question

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lxl +x=4-3x just want to know how to set it up this a total new topic and pretty confusing and got a quiz and stuff on it friday so want to get it in my head you know so would mean alot.

OpenStudy (anonymous):

Okay first I want you to get my math tutoring app in Itunes or android if you have a device, type Trevor Doyle in the app store. They are videos of me teaching but here we go:

OpenStudy (anonymous):

its free?

OpenStudy (anonymous):

I don't mean you mean you need to get for me to help

OpenStudy (anonymous):

no sorry, but really cheap

OpenStudy (anonymous):

hm im purchasing an itunes card soon so ill def check it out!

OpenStudy (anonymous):

okay here we go

OpenStudy (anonymous):

theres one for free I see

OpenStudy (anonymous):

remember you need to get this: absolute value = everything else

OpenStudy (anonymous):

yes but you need teh algebra 1

OpenStudy (anonymous):

im def going to download it if end up getting the card but thanks for letting me know about it looks legit.

OpenStudy (anonymous):

lxl +x=4-3x -x -x

OpenStudy (anonymous):

|x| =4-4x

OpenStudy (anonymous):

then whatever is in the | |, write it twice the make one equal to positive (4-4x) and the other negative x = +(4-4x) x = -(4-4x)

OpenStudy (anonymous):

and distrib

OpenStudy (anonymous):

x = 4-4x and x=-4+4x

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

got it thanks!

OpenStudy (anonymous):

coolness

OpenStudy (anonymous):

Respect ^^

OpenStudy (anonymous):

:)

hero (hero):

There's plenty of FREE math tutoring videos out there

OpenStudy (anonymous):

yes there are but not for the iphone

hero (hero):

You can access the internet with iphone, yes?

hero (hero):

You can access youtube with an iphone, yes?

OpenStudy (anonymous):

that's fine, you get what you pay for, I am Math teacher with a Masters Degree, but you can have the free one if you want, good luck?

hero (hero):

How about khan academy with an iphone? yes

OpenStudy (anonymous):

So use that

hero (hero):

You get what you pay for.....You're thinking about money too much I believe.

OpenStudy (anonymous):

I have techniques that really work, and I have helped hundreds so I am worth it, and its only $4.99 for 100 videos

hero (hero):

Good for you, but I think it would have been better if you posted a donate button like Khan and Patrick

OpenStudy (anonymous):

searchable, organized together

OpenStudy (anonymous):

I am doing fine, 100's of downloads monthly

OpenStudy (anonymous):

13000 in one month

OpenStudy (anonymous):

http://itunes.apple.com/us/app/algebra./id496511723?ls=1&mt=8 app store

OpenStudy (anonymous):

android: https://play.google.com/store/apps/developer?id=Trevor+Doyle

hero (hero):

I was just helping you out. Your name is not exactly searchable via app store

OpenStudy (anonymous):

your search was in the apple store, my app is searchable in iTunes app store and android store, but thanks